A wave has a length of 2m and a frequency of 10Hz. Calculate the wave number . Calculate the angular frequency. Calculate wave speed


1
Expert's answer
2021-07-01T15:29:50-0400

The wave number:

"k=\frac{2 \pi}{λ} \\

= \frac{2 \times 3.14}{2}= 3.14 \;m^{-1}"

The angular frequency:

"ω= \frac{2 \pi}{T}"

T=the period of the wave

"T=\frac{1}{f} \\

= \frac{1}{10}=0.1"

"ω= \frac{2 \times 3.14}{0.1}=62.8 \;s^{-1}"

Wave speed:

"v= \frac{ω}{k} \\

= \frac{62.8}{3.14}=20 \;m/s"